The perimeter of a triangle is 43 inches. The second side is three inches longer than the first side. The third side is one inch longer than the second. Find the length of each side.
Answer by wuwei96815(245)

Let x=first side, x+3=second side, and x+3+1=third side.

All together they should total 43. Therefore:
(x)+(x+3)+(x+3+1) = 43

Remove the parentheses and solve for x.
x+x+3+x+3+1 = 43
3x + 7 = 43
3x = 43-7 = 36
x = 36/3 = 12

x+3 = 12+3 = 15
x+3+1 = 12+3+1 = 16